OpenStudy (anonymous):

Molly finds two sticks and wants to build a kite for her little brother. The sticks are 16 inches and 38 inches long, and she plans on using them as the diagonals of the kite. Molly will need a minimum of _____ square inches of fabric to make the kite. *This is confusing b/c i got a number with a decimal and 1st of all idk if i am right, 2nd of all should i round it?* i got 82.462112

OpenStudy (anonymous):

well why is she going to make one you know they sell them at walmart

OpenStudy (anonymous):

just being silly(:

OpenStudy (anonymous):

what?

OpenStudy (ajprincess):

Is it clear?

OpenStudy (anonymous):

yeah. thanks :)

ganeshie8 (ganeshie8):

how you got 82.462112 ?

OpenStudy (anonymous):

@ganeshie8 i realized what i did it wrong

OpenStudy (ajprincess):

I am sorry.  Area of the triangle=1/2*16*19 Area of the rhombus=2*1/2*16*19 =16*19 =304 It was nt mod dr. it was me who deleted t. i made a mistake in t.